OpenStudy (anonymous):

In the problem of "e^(x)-e^(-x)=2", how do you get to here, "e^(2x)-2e^(x)-1=0".

jhonyy9 (jhonyy9):

us ln on both sides

jhonyy9 (jhonyy9):

2 will be 2lne what signe ln(e)2 logarithm natural from e on exponent 2

OpenStudy (ash2326):

just convert e^-x as 1/e^x we have e^x-1/e^x=2 multiplying whole equation by e^x we get e^2x-1=2e^x so we get e^2x-2e^x=1

OpenStudy (anonymous):

blah thats what I did I just messed up the last step thanks :D

jhonyy9 (jhonyy9):

so but if you need to solve it you need to us what i have wrote before

OpenStudy (anonymous):

No to solve it you need to use the quadratic formula

OpenStudy (jamesj):

CP is exactly right.